### v2

@@ 3.3 Baseline Battery Controllers
The battery management system was optimized through the implementation of two rule-based strategies which are MSC and TOU in the baseline algorithm b5 . The MSC is a means of optimizing the utilization of surplus energy generated by the PV system through its storage in a battery. The TOU involved modifying the battery charging process in response to variations in electricity prices. These two strategies were implemented to manage the battery as a baseline comparison method.
@@ 3.4 Q Learning
Q-learning algorithms employ the Bellman equation b_bellman to choose maximum Q-values and the generalized Bellman equation is expressed in Equation 4.
@@
Equation 4 presents the correlation between the action-value function $Q_{\pi}$, the reward, and transition probabilities of the environment. The statement specifies that the value of $Q_{\pi}(s,a)$ is equivalent to the summation of the probability $p(s^{\prime},r|s,a)$ of transitioning to state ${}^{\prime}$ and receiving reward R, multiplied by the summation of the immediate reward R and the discounted value of the subsequent state ${}^{\prime}$ under the policy $\pi$, considering all feasible next states ${}^{\prime}$ and rewards R. The parameter $\gamma$, commonly referred to as the discount factor, plays a crucial role in determining the relative significance of rewards that are obtained immediately versus those that are obtained in the future. The Bellman equation is a fundamental concept within the field of RL, used used for numerous algorithms that aim to acquire knowledge regarding the value function and policy optimization.
Q-learning involves using the current estimate of $Q^{\pi}$ to improve its future predictions by including the known reward value $r(s_{t},a_{t})$. Q-learning fundamentally relies on the concept of Temporal Difference (TD) learning b11 . In this method, the Q-value is updated after performing an action in the state $S_{t}$ and observing the resulting reward $r_{t}$ which leads to a transition to the next state $s_{t+1}$. The TD is mathematically represented in the Equation 2.
Empirical evidence supports the notion that as the frequency of visits to each state-action pair’s Q-value approaches infinity, the learning rate $\alpha$ exhibits a decreasing trend concerning the time step t. As the value of t approaches infinity, the function $Q(s;a)$ approaches the optimal $Q*(s;a)$ for all possible state-action pairs b12 . In this study, the Q-learning algorithm was utilized to optimize the management of battery charging and discharging operations to reduce the cost of imported electricity from the power grid. The Q-learning algorithm comprises different components, namely the state space denoted as $S$, the action space represented by $A$, and the reward function, which is the aggregate cost of electricity denoted as $R$.
@@ 3.5 Application of Q-Learning to Battery Management
In this study, Q-learning is employed as a means of effectively managing the process of battery charging and discharging. This is achieved through the exploration of the state space and action space, which are integral components of the environment. The reward is calculated by considering various actions, such as charging, discharging, or remaining idle, in response to factors such as renewable generation and electricity prices. The state space, action space, and reward are explained below. The proposed algorithm is illustrated in the flow chart shown in Figure 4. The algorithm begins by initializing the environment, specifying the available actions, defining a strategy for computing rewards, and determining the number of episodes. Subsequently, it initialize the learning rate and exploration rate to 0.8, the discount factor to 0.9, and initializes the Q Table to 0. The algorithm employs the weight decay with the decay of 0.0001 to gradually decrease the learning rate and exploration rate with respect to the episodes. To determine the appropriate action, the algorithm uses the epsilon-greedy policy.
@@ 4.1 Q-Learning for Battery management
In this scenario, the Q-learning algorithm was trained to enhance the efficiency of battery management in dairy farming. Its primary goal is to increase the use of PV energy while reducing dependence on the external power grid and to lower energy cost in the dairy farm. This algorithm was trained on one year’s data from Finland b10 . The trained algorithm learned the optimal policy for charging the battery, discharging, and remaining idle, considering state information on battery charge level, time, and energy prices. After training, the algorithm’s performance was tested on the same dataset for one year. The findings indicate that the implementation of the Q-learning algorithm decreases the import of electricity by 10.64%. In comparison, the baseline strategy resulted in a decrease in electricity imports only by 9.72%. This improvement in the reduction of electricity imports from the grid is presented in Figure 5, demonstrates the algorithm’s effectiveness.
@@
Figure 6 presents a side-by-side comparison of battery charge behaviors over two days, using two distinct methodologies: baseline and Q-learning. The x-axis of the graph indicates the hours of the day, while the y-axis demonstrates the battery’s charge in kWh. This comparison shows the behavior of the battery charging and discharging using different methodologies. The Q-learning method demonstrates a more effective approach to battery management, considering various factors: PV generation, electricity prices, and the energy demands of the farm. It charges the battery when electricity prices are low and PV generation is available, ensuring efficient use of PV generation. During peak hours and periods of high electricity prices, the Q-learning strategy discharges the battery. In contrast, the TOU method follows a more static approach, based on predetermined rules while Q-learning is adaptive to the current environment. This adaptability allows for more effective management of battery charging and discharging, aligning with fluctuating energy demands and variable PV generation, leading to enhanced efficiency and cost savings for the dairy farm.
@@ 4.3 Investigating State Space
The Q-learning algorithm is trained and tested with different state spaces including scenarios with and without wind generation data. Figure 9 compares the algorithm’s performance across these different state spaces. Figure 8(a) shows how the inclusion of load and PV generation in the state space affects electricity import reduction, compared to the state space from experiment 4.1. We found that the state space from experiment 4.1 is more effective, reducing electricity imports by 10.64%, compared to the modified state space (with load, and PV) which only achieved a 9.97% reduction. Figure 8(b) illustrates the impact of incorporating wind generation data into the state space. When the state space from experiment 1 is combined with wind data, there is a significant reduction in load import, achieving a 22.14% decrease. In contrast, a modified state space that includes load, PV, and wind data results in a smaller reduction of only 10.07%. This shows that expanding the state space adds to the complexity of the environment, which makes it difficult for the Q-learning agent to make optimal decisions. Another reason for this incapability could be due to the curse of dimensionality in Q-learning, where increasing dimensionality leads to sparser data and challenges in achieving expected resultsb11 . Additionally, discretizing the state space to manage its dimensionality might have affected performance due to variations in the data.
@@ 5 Conclusion
In this research, the Q-learning is applied to battery management in a dairy farm, using electricity data from Finland. This study involved various experiments to assess the effectiveness of the Q-learning algorithm. This research explored the effect of integrating wind and solar data on battery management and examined how changing the state space of the algorithm impacts its performance. Additional experiments were conducted using data from Ireland to validate the algorithm’s effectiveness. As explained in Section 4, findings show that the Q-learning algorithm successfully reduced the dairy farm’s reliance on the external grid.

### v3

@@ 3.3 Baseline Battery Controllers
The battery management system was optimized through the implementation of two rule-based strategies which are MSC and TOU in the baseline algorithm [8]. The MSC is a means of optimizing the utilization of surplus energy generated by the PV system through its storage in a battery. The TOU involved modifying the battery charging process in response to variations in electricity prices. These two strategies were implemented to manage the battery as a baseline comparison method.
@@ 3.4 Q Learning
Q-learning algorithms employ the Bellman equation [48] to choose maximum Q-values and the generalized Bellman equation is expressed in Equation 4.
@@
Equation 4 presents the correlation between the action-value function $Q_{\pi}$, the reward, and transition probabilities of the environment. The statement specifies that the value of $Q_{\pi}(s,a)$ is equivalent to the summation of the probability $p(s^{\prime},r|s,a)$ of transitioning to state S and receiving reward R, multiplied by the summation of the immediate reward R and the discounted value of the subsequent state S under the policy $\pi$, considering all feasible next states S and rewards R. The parameter $\gamma$, commonly referred to as the discount factor, plays a crucial role in determining the relative significance of rewards that are obtained immediately versus those that are obtained in the future. The Bellman equation is a fundamental concept within the field of RL, used used for numerous algorithms that aim to acquire knowledge regarding the value function and policy optimization.
Q-learning involves using the current estimate of $Q^{\pi}$ to improve its future predictions by including the known reward value $r(s_{t},a_{t})$. Q-learning fundamentally relies on the concept of Temporal Difference (TD) learning [49]. In this method, the Q-value is updated after performing an action in the state $S_{t}$ and observing the resulting reward $r_{t}$ which leads to a transition to the next state $s_{t+1}$. The TD is mathematically represented in the Equation 2.
Empirical evidence supports the notion that as the frequency of visits to each state-action pair’s Q-value approaches infinity, the learning rate $\alpha$ exhibits a decreasing trend concerning the time step t. As the value of t approaches infinity, the function $Q(s;a)$ approaches the optimal $Q*(s;a)$ for all possible state-action pairs [11]. In this study, the Q-learning algorithm was utilized to optimize the management of battery charging and discharging operations to reduce the cost of imported electricity from the power grid. The Q-learning algorithm comprises different components, namely the state space denoted as $S$, the action space represented by $A$, and the reward function, which is the aggregate cost of electricity denoted as $R$.
@@ 3.5 Application of Q-Learning to Battery Management
In this study, Q-learning is employed as a means of effectively managing the process of battery charging and discharging. This is achieved through the exploration of the state space and action space, which are integral components of the environment. The reward is calculated by considering various actions, such as charging, discharging, or remaining idle, in response to factors such as renewable generation and electricity prices. The state space, action space, and reward are explained below. The proposed algorithm is illustrated in the flow chart shown in Figure 4. The algorithm begins by initializing the environment, specifying the available actions, defining a strategy for computing rewards, and determining the number of episodes. Subsequently, it initializes the learning rate and exploration rate to 0.8, the discount factor to 0.9, and initializes the Q Table to 0. The algorithm employs the weight decay with the decay of 0.0001 to gradually decrease the learning rate and exploration rate concerning the episodes. To determine the appropriate action, the algorithm uses the epsilon-greedy policy.
@@ 4.1 Q-Learning for Battery management
In this scenario, the Q-learning algorithm was trained to enhance the efficiency of battery management in dairy farming. Its primary goal is to increase the use of PV energy while reducing dependence on the external power grid and to lower energy cost in the dairy farm. This algorithm was trained on one year’s data from Finland [42]. The trained algorithm learned the optimal policy for charging the battery, discharging, and remaining idle, considering state information on battery charge level, time, and energy prices. After training, the algorithm’s performance was tested on the same dataset for one year. The findings indicate that the implementation of the Q-learning algorithm decreases the import of electricity by 10.64%. In comparison, the baseline strategy resulted in a decrease in electricity imports only by 9.72%. This improvement in the reduction of electricity imports from the grid is presented in Figure 5, demonstrates the algorithm’s effectiveness.
@@
Figure 6 illustrates a comparison of battery charging behaviors throughout a day using two methodologies: baseline and Q-learning. Additionally, it displays electricity price, consumption, PV, and wind generation data for the first day of the year. The x-axis represents the hours of the day, while the y-axis indicates the battery and electricity profiles on the farm. This comparison highlights differences in battery charging and discharging behaviors between the two methodologies. The Q-learning method demonstrates enhanced battery management, with results indicating an optimal policy for charging and discharging. Specifically, when it charges the battery during periods of low electricity prices and available PV and wind generation, maximizing the utilization of renewable energy sources. Conversely, during peak hours when electricity prices are high, the Q-learning algorithm discharges the battery. In contrast, the rule-based method follows a more static approach, based on predetermined rules while Q-learning is adaptive to the current environment. This adaptability allows for more effective optimization of the battery charging and discharging, aligning with fluctuating energy demands and variable PV and wind generation, leading to enhanced efficiency and cost savings for the dairy farm.
@@ 4.3 Investigating State Space
The Q-learning algorithm is trained and tested with different state spaces including scenarios with and without wind generation data. Figure 9 compares the algorithm’s performance across these different state spaces. Figure 9(a) shows how the inclusion of load and PV generation in the state space affects electricity import reduction, compared to the state space from experiment 4.1. We found that the state space from experiment 4.1 is more effective, reducing electricity imports by 10.64%, compared to the modified state space (with load, and PV) which only achieved a 9.97% reduction. Figure 9(b) illustrates the impact of incorporating wind generation data into the state space. When the state space from experiment 1 is combined with wind data, there is a significant reduction in load import, achieving a 22.14% decrease. In contrast, a modified state space that includes load, PV, and wind data results in a smaller reduction of only 10.07%. This shows that expanding the state space adds to the complexity of the environment, which makes it difficult for the Q-learning agent to make optimal decisions. Another reason for this incapability could be due to the curse of dimensionality in Q-learning, where increasing dimensionality leads to sparser data and challenges in achieving expected results[49]. Additionally, discretizing the state space to manage its dimensionality might have affected performance due to variations in the data.
@@ 5 Conclusion
In this research, Q-learning is applied to battery management in a dairy farm, using electricity data from Finland. This study involved various experiments to assess the effectiveness of the Q-learning algorithm. This research explored the effect of integrating wind and solar data on battery management and examined how changing the state space of the algorithm impacts its performance. Additional experiments were conducted using data from Ireland to validate the effectiveness of the algorithm. As explained in Section 4, the findings show that the Q-learning algorithm successfully reduced the reliance of the dairy farm on the external grid.